OpenStudy (anonymous):

determine the acute angle between the line 2x+3y=6 medal provided

OpenStudy (anonymous):

between what ?

OpenStudy (anonymous):

and the y axis

OpenStudy (anonymous):

 in y=mx+c form m is slope of line m=2 tan(P)=-2/3 angle \[P=\tan^{-1} (-2/3) \] Q=180-P R=90-Q=90-180+P=P-90